OpenStudy (anonymous):

Fifteen more than four times a number is equal to the difference between 50 and three times the number. Find the number.

OpenStudy (anonymous):

let the number = x four times the number = 4x three times the number = 3x ATQ, 4x + 15 = 50 - 3x 4x + 3x = 50 - 15 7x = 35 x = 35/7 x = 5

OpenStudy (anonymous):

so the number is 5